OpenStudy (anonymous):

Can any show me how to find the angle between the two vectors (-7,1,1) and (3,2,-2)

OpenStudy (amistre64):

you will need 3 things; at least .... the dot product, and the length of each vector

OpenStudy (amistre64):

\[cos(\alpha)=\frac{a\cdot b}{|a|~|b|}\] \[\alpha=cos^{-1}\left(\frac{a\cdot b}{|a|~|b|}\right)\]